OpenStudy (surana):

If y varies directly as x, find the constant of variation if y = 12 when x = 1/3.

OpenStudy (surana):

The possible answers are... A: 36 B: 4 C: 6 D: 12

OpenStudy (dtan5457):

direct variation is made by the formula of y=kx plug that in, solve for k.

OpenStudy (surana):

So technically, it's 12 = 36/4/6/12 * 1/3?

OpenStudy (dtan5457):

i don't really know what you mean but 12=k(1/3) what times 1/3 =12

OpenStudy (surana):

What I meant was that I put 12 as Y, times K, which could be either 36, 4, 6 or 12, multiplied by 1/3. Is that what I do? Or do I do something else?

OpenStudy (dtan5457):

i think you are making is harder than it has to be 12=k(1/3) whatever answer choice you have satisfies k that makes that ^ equation true, is your constant.

OpenStudy (dtan5457):

which means there is only one answer

OpenStudy (surana):

Oh, okay. So I divide 12 by 1/3?

OpenStudy (dtan5457):

yes, that would work.

OpenStudy (surana):

Okay, give me a minute or so.

OpenStudy (dtan5457):

np

OpenStudy (surana):

I got 4. But I don't think that's right.

OpenStudy (dtan5457):

if your dividing by less than 1, the number always turns out greater than the original number

OpenStudy (dtan5457):

check that you put parenthesis.

OpenStudy (surana):

So it should be different, I'll try again.

OpenStudy (surana):

Wait, is this a flip and multiply thing? Because if it is then the answer should be 36.

OpenStudy (dtan5457):

correct.

OpenStudy (surana):

So 36 is the answer by the calculations we did?

OpenStudy (dtan5457):

yep

OpenStudy (surana):

Okay, just to check, Y = 12, x = 1/3, then that becomes 12/1 * 3/1, which becomes 36/1 and thus 36, am I correct?

OpenStudy (dtan5457):

yes

OpenStudy (surana):

That's good, I'll wait to see what triciaal has to say first, though.

OpenStudy (triciaal):